OpenStudy (anonymous):

Write the equation of the line that passes through (0, 3) and is parallel to the line 3x + 5y = 6.

OpenStudy (anonymous):

do you know what the slope is in 3x + 5y = 6 ? If not, do you know how to find it ?

OpenStudy (anonymous):

In y = mx + b form, the m position is the slope 3x + 5y = 6 (put this in y = mx + b form) 5y = -3x + 6 y = -3/5x + 6/5 the slope(m) = -3/5 A parallel line will have the same slope. now we use y = mx + b slope(m) = -3/5 (0,3) x = 0 and y = 3 now sub y = mx + b 3 = -3/5(0) + b 3 = b your equation is : y = -3/5x + b...but we need it in standard form Ax + By = C y = -3/5x + 3 -- multiply by 5 to get rid of the fraction 5y = -3x + 3 -- add 3x to both sides 3x + 5y = 3 <== your parallel line